### v1

# RLGNet: Repeating-Local-Global History Network for Temporal Knowledge Graph Reasoning

Abstract
Temporal Knowledge Graph (TKG) reasoning is based on historical information to predict the future. Therefore, parsing and mining historical information is key to predicting the future. Most existing methods fail to concurrently address and comprehend historical information from both global and local perspectives. Neglecting the global view might result in overlooking macroscopic trends and patterns, while ignoring the local view can lead to missing critical detailed information. Additionally, some methods do not focus on learning from high-frequency repeating events, which means they may not fully grasp frequently occurring historical events. To this end, we propose the Repetitive-Local-Global History Network(RLGNet). We utilize a global history encoder to capture the overarching nature of historical information. Subsequently, the local history encoder provides information related to the query timestamp. Finally, we employ the repeating history encoder to identify and learn from frequently occurring historical events. In the evaluation on six benchmark datasets, our approach generally outperforms existing TKG reasoning models in multi-step and single-step reasoning tasks.
1 Introduction
TKG is a structured yet highly complex knowledge system. In TKG, each fact is represented by a quadruple, including subject, relation, object, and timestamp, thereby providing detailed and accurate temporal information. TKG allow us to understand the relationships between entities and capture the changes in these relationships via timestamps.
TKG extrapolation aims to predict new facts at a future timestep based on the ground truth facts within the known timeframe. Extrapolation is further classified into two types: (1) Single-step reasoning predicts future events using known ground truth facts from past timestamps within the test period. (2) Multi-step reasoning forecasts without relying on ground truth facts from the test period. By analyzing past events and identifying factors that influence event progression, we can infer the probabilities of emergent future facts. However, since future events contain unknown factors, TKG extrapolation is a challenging problem(?).
Figure 1: A figure to explain repeating, local, and global facts. ”Frequency” represents the number of occurrences of that type of event, ”Interval” represents the time difference from the last occurrence of that type of event to the time of query. The red arrows represent the facts most relevant to the query.
Predicting the future is a task that requires a deep analysis of historical data, the key to which is to discern possible patterns and trends from it. In previous work, it’s generally believed that the facts closer to the query time are more important. However, this method may cause us to overlook some facts that have the potential to impact future trends, because these facts are not necessarily always close to the query time. As shown in Figure 1, (United States, Negotiations, ?) does not appear in the local history. According to historic recurrence(?), historical facts could have repeated or even cyclical patterns. As illustrated in the same Figure 1, (United States, Negotiations, North Korea) has occurred 95 times. Therefore, we need to consider the repetitiveness, locality, and globality of historical information. Moreover, some work has shown that not every fact in the TKG is crucially important for prediction. To improve the accuracy of prediction, we need to selectively pay attention to the facts that are most relevant to the query, as the fact indicated by the red arrow in Figure 1. Therefore, finding a method that is based on the query and simultaneously considers repeating, local and global historical information is of paramount importance for extrapolation in TKG.
To address this challenge, we design and propose a model called Repeating-Local-Global History Network (RLGNet). In this model, we design repeating, local, and global history encoders respectively based on the query-oriented method. In the local history encoder, the model views the TKG as a Knowledge Graph (KG) sequence and models the facts in adjacent timestamps to obtain local historical information. In the global history encoder, the model aggregates all facts using an attention mechanism to gather global historical information. Finally, building on the foundation of local and global historical encoders, we use a repeating history encoder to further increase the scores of repetitive events. Overall, this paper makes the following contributions:
  1. 1.
    We designed three different levels of encoders to deal with the repetitive, local, and global historical information respectively.
  2. 2.
    We explored the contributions of repeating, global and local historical information to predictive performance in two different extrapolation settings.
  3. 3.
    We conducted extensive experiments on six public TKG datasets, proving the model’s effectiveness in entity prediction.
2 Related Work
Static KG Reasoning: Recent years have seen great interest and research in static Knowledge Graph (KG) reasoning models.(??) These models include distance-based models, such as TransE(?) and TransH(?), which determine the likelihood of facts by measuring the distances between entities. Another category is semantic-matching-based models, like DistMult(?) and RESCAL(?). There are also models based on Convolutional Neural Networks (CNNs), such as ConvE(?) and Conv-TransE(?), which represent entities and relations using matrices processed by convolutional kernels. Similarly, Graph Convolutional Network (GCN) models, including R-GCN(?) and VR-GCN(?), stand out for their capacity to integrate graph structures with node features. However, these models focus on static KG, and their predictive capabilities for future events are limited.
Temporal KG Reasoning: TKG reasoning has two settings: extrapolation and interpolation. The interpolation setting aims at predicting missing historical facts rather than future events. In contrast, this paper focuses on reasoning within the extrapolation setting to forecast future facts using historical data. Know-Evolve(?) and DyREP(?) model the occurrence of facts in TKG using a temporal point process. Glean(?) enriches factual features by utilizing unstructured text information. CyGNet(?) captures and understands historical trends and patterns through a replicating generative mechanism. CENET(?) distinguishes events as historical and non-historical, using this classification for contrastive learning. Some models, such as RE-GCN(?), utilize Graph Convolutional Networks (GCNs) to simulate the evolutionary process of knowledge graphs, thereby capturing and learning the dynamic properties of entities and relations. TiRGN(?) uses a local-global historical approach for reasoning. CEN(?) and CluSTeR(?) utilize reinforcement learning to capture and learn the connections between entities and relationships in the knowledge graph. TLogic(?) utilizes temporal logic rules to constrain the predicted paths for queries. TANGO(?) utilizes neural ordinary differential equations to model the structural information of each entity. However, these models often focus on partial aspects of historical information without combining query data with historical details. Consequently, they can’t accurately capture recurring, local, and global facts.
3 Preliminaries
TKG represent events temporally as snapshots. Let $\mathcal{G}_{raw}$ be the sequence of snapshot graphs $\mathcal{G}_{raw}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}\}$. Each fact in the snapshot is a quadruple $(s,r,o,t)$, where $s,o\in\mathcal{E},r\in\mathcal{R},t\in\mathcal{T}$. It represents a relationship $r$ between subject entity $s$ and object entity $o$ at time $t$. For each quadruple $(s,r,o,t)$,an inverse relation quadruple$(o,r^{-1},s,t)$ is often added to the dataset. The entity prediction task aims to predict the missing entity in a query $q=(s_{q},r_{q},?,t_{q})$. We denote all entity embeddings in dimension $d$ with $H\in\mathbb{R}^{|\mathcal{E}|\times d}$, and relation embeddings similarly with $R\in\mathbb{R}^{|\mathcal{R}|\times d}$. In each encoder, the initial weights of H and R are randomly generated. The specific entity and relation are represented as $h\in\mathbb{R}^{d}$and $r\in\mathbb{R}^{d},$ respectively.
Given the query tuple $q=(s_{q},r_{q},?,t_{q})$, we define the candidate entity set at time $t$ as $\mathcal{C}_{q}^{t}=\{o|(s_{q},o_{q},o,t)\in\mathcal{G}_{t}\}$. The candidate entity set over a time range $t_{1}$ to $t_{2}$ is defined as $\mathcal{C}_{q}^{t_{1}:t_{2}}$, a union of all candidate sets over the range.
$\mathcal{C}_{q}^{t_{1}:t_{2}}=\bigcup_{i=t_{1}}^{t_{2}}\mathcal{C}_{q}^{i}$ (1)
Since not all candidate entities greatly affect the prediction, we only keep the top $k$ most frequent entities, denoted as $C_{q,top_{k}}^{1:t_{q}-1}$. Candidates in the set usually represent repeating facts, we construct $\mathcal{G}_{rep}$, containing all tuples in the period where o is a top $k$ candidate before $t-1$. $\mathcal{G}_{rep}$ can be formalized as:
$\mathcal{G}_{rep}=\{q=(s_{q},r_{q},o_{q},t_{q})\in\mathcal{G}_{raw}|o\in\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}\}$ (2)
$\mathcal{G}_{rep}$ essentially extracts prominent, repetitive events from historical data.
Figure 2: The overall architecture of RLGNet. The upper half is the local history encoder which learns local historical information through KG sequences of adjacent timestamps. The lower half comprises the repeating and global history encoders, learning repetitive and global historical information respectively by statistically querying candidate entities. The content within the dashed box represents the entity most relevant to the query. The left side sums the scores of the three parts to get the final prediction score.
4 Model Overview
The overall framework of the model is illustrated in the Figure 2. This model can be divided into three core submodules: the local history encoder, the global history encoder and the repeating history encoder. The local history encoder captures local facts by exploring the structural features and historical dependencies the sequence of KG at adjacent timestamps. The global history encoder captures global facts by examining the relevant facts across all previous timestamps and capturing entities and relationships that may not be evident at adjacent timestamps. Furthermore, these two encoders use a scoring decoder to evaluate the facts. The primary task of the repeating history encoder is to enhance its ability to predict repetitive facts by encoding frequently occurring historical events. All encoders utilize time vector to encode time and frequency. By utilizing these three encoders, the model is capable of making accurate predictions in response to queries.
4.1 Scoring Decoder
Inspired by the work of ConvE, we designed this scoring decoder and used it in the next two decoders. First, the model reshapes a one-dimensional input into a two-dimensional matrix. Then, it processes this matrix with a CNN. After that, it flattens the resulting data and outputs the embeddings for the target entity through a fully connected layer.
$\varphi(H,V_{1}\dots V_{k})=\sigma(ConvE(V_{1},\dots,V_{k}))\cdot H$ (3)
$\sigma$ represents the ReLu function. $V_{i}(1\leq i\leq k)$are input vectors, $\varphi(H,V_{1},\dots,V_{k})\in\mathbb{R}^{|\mathcal{E}|}$ represents predicted entity scores.
4.2 Time Vector
In terms of time, the module employs two types of temporal encoding vectors, periodic(?) and aperiodic, to represent time.
$\textbf{v}_{1}(t)&=\cos(W_{1}^{t}t+b_{1})\\ \textbf{v}_{2}(t)&=\tanh(W_{2}^{t}t+b_{2})\end{aligned}\right.$ (4)
Then we concatenate the two vectors to obtain the final vector representation.
$\textbf{V}(t)=[\textbf{v}_{1}(t);\textbf{v}_{2}(t)]$ (5)
$W_{1}^{t},W_{2}^{t},b_{1},b_{2}$ are learnable parameters and the vectors $\textbf{v}_{1}(t)$ and $\textbf{v}_{2}(t)$ have the same dimensionality, [;] denotes the vector concatenation operation, $\textbf{V}(t)$ is the time vector.
4.3 Local History Encoder
This module captures local facts by focusing on adjacent history. For each query $q=(s_{q},r_{q},?,t_{q})$, the module pays attention to the $m$ timestamp subgraph $\{\mathcal{G}_{t_{q}-m},\dots,\mathcal{G}_{t_{q}-1}\}$ related to this query to obtain the structural features of the subgraph. For this purpose, we employ GCN to aggregate individual subgraphs, use Gated Recurrent Units (GRU)(?) to learn the features of subgraph evolution, and adopt an attention mechanism to integrate information from candidate entities. Firstly, we use GCN with $\omega$ layers to get the entity representations at the current moment.
$h_{o}^{t,l+1}=\sum\limits_{(s,r,o)\in\mathcal{G}_{t}}\frac{1}{|N_{o}^{t}|}W_{1}^{l}\phi(h_{i}^{t,l},r_{t})+W_{2}^{l}h_{o}^{l}$ (6)
$\phi(h_{i}^{t,l},r_{t})=[h_{i}^{t,l};r_{t};h_{i}^{t,l}+r_{t};h_{i}^{t,l}\cdot r_{t}]$ (7)
The symbol [$\cdot$] represents the Hadamard product. $N_{o}^{t}$ represents neighbors of node $o$ at time $t$. The embedded representation of node $o$ at the $l$-th layer is denoted as $h_{o}^{t,l}\in\mathbb{R}^{d}$. The aggregation and self-loop parameters of the $l$-th layer are shown as $W_{1}^{l}$ and $W_{2}^{l}$. The edge relation type is presented as $r_{t}$. Specifically, when $\omega=0$, i.e., $h_{o}^{t,0}=\sum_{i\in N_{o}^{t}}\frac{1}{|N_{o}^{t}|}h_{i}^{t}$. Entity representations are combined with their temporal representations and then predicted for the next moment with GRU:
$H_{t+1}=GRU([H_{t};\textbf{V}_{1}(t_{q}-t)],H^{{}^{\prime}})$ (8)
$\textbf{V}_{1}(t_{q}-t)\in\mathbb{R}^{d_{2}}$ is the time vector. $H^{{}^{\prime}}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the final GCN layer output. The attention mechanism aggregates candidate entities over a history of length $m$. Node aggregation for query $q$ is:
$C_{q}^{t}=\sum\limits_{i\in C_{q}^{t}}\frac{1}{|C_{q}^{t}|}h_{i}^{{}^{\prime}}$ (9)
$C_{q}^{t}$ represents query $q$’s candidate entity representation at time $t$. Next, the module calculates attention weights $a_{q}^{t}$ for the query at time $t$:
$a_{q}^{t}=W_{2}\sigma(W_{1}[h_{q}^{{}^{\prime}};r_{q};\textbf{V}_{1}(t_{q}-t);C_{q}^{t}])$ (10)
$W_{1}$, and $W_{2}$ are learnable parameters. After calculating attention weights, these are used to aggregate candidate entities for query q:
$C_{q}^{{}^{\prime}}=\sum_{i=1}^{m}\frac{\exp{(a_{q}^{i})}\cdot C_{q}^{i}}{\sum_{j=1}^{m}\exp{(a_{q}^{j})}}$ (11)
The candidate entity representation $C_{q}^{{}^{\prime}}$ for query $q$. The module then calculates the entity score for query $q$’s local history encoder:
$score_{loc}^{q}=\varphi_{loc}(H^{{}^{\prime}},h_{q}^{{}^{\prime}},r_{q},C_{q}^{{}^{\prime}})$ (12)
Eq (12) can be viewed as an instance of Eq (3).
4.4 Global History Encoder
The global history encoder is designed to consider and extract information from a global perspective in order to capture global facts. For each query, we count the frequencies and the recent appearances of candidate entities. The module then uses an attention mechanism to integrate this information about the candidate entities. First, the attention scores for the time-related candidate entities of the current query are calculated:
$\mathcal{A}_{1}(q,i)=\sigma(W_{3}[h_{q};r_{q}])^{T}\sigma(W_{4}[\textbf{V}_{2}(t_{q}-t_{q}^{i});h_{i}])$ (13)
The attention scores for frequency-related candidates is similarly:
$\mathcal{A}_{2}(q,i)=\sigma(W_{5}[h_{q};r_{q}])^{T}\sigma(W_{6}[\textbf{V}_{3}(cnt^{i}_{q});h_{i}])$ (14)
Considering the similarities between the characteristics of frequency and time, we employ a time vector to represent frequency. $W_{4}\in\mathbb{R}^{d\times 2d},W_{3}\in\mathbb{R}^{d\times 2d},W_{5}\in\mathbb{R}^{d\times 2d}$, and $W_{6}\in\mathbb{R}^{d\times 2d}$ are learnable parameters. $t_{q}^{i}$ and $cnt_{q}^{i}$ represent the frequency and last occurrence of candidate entity $i$ for query q. Additionally, taking computational resources into consideration, we only selected $top_{k}^{all}$ candidate entities. Following, we normalize their respective scores and multiply by the candidate entity representations to obtain representations $C_{q}^{gap}$ and $C_{q}^{cnt}$ for time-relevant and frequency-relevant candidates, respectively.
$C_{q}^{gap}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{1}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{1}(q,j))}$ (15)
$C_{q}^{cnt}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{2}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{2}(q,j))}$ (16)
The module then calculates the entity score for the global history encoder:
$score_{glo}^{q}=\varphi_{glo}(H,h_{q},r_{q},\textbf{V}_{4}(t_{q});C_{q}^{gap};C_{q}^{cnt})$ (17)
$h_{q}$ and $r_{q}$ represent the query entity and relation, Eq (17) can be viewed as an instance of Eq (3).
4.5 Repeating History Encoder
The Repeating History Encoder is designed to improve prediction by increasing the weight of historical events. We first filter out and retain all the repetitive facts. Then, only the candidate entities of facts that rank in the top($top_{k}$) in terms of frequency are retained. In this module, the frequency of the candidate entities is also represented using a time vector. The repeating history encoder’s score function is a multilayer perceptron (MLP):
$\textbf{MLP}(x)=W_{7}\cdot\sigma(W_{8}\cdot\sigma(W_{9}\cdot x))$ (18)
$W_{7}\in\mathbb{R}^{1\times d}$, $W_{8}\in\mathbb{R}^{d\times 3d}$, and $W_{9}\in\mathbb{R}^{3d\times 5d}$ are learnable parameters. The extra score for candidates is calculated as follows:
$score_{rep}^{q,i}=\textbf{MLP}([h_{q};r_{q};\textbf{V}_{5}(t_{q});h_{i};\textbf{V}_{6}(cnt_{q}^{i})])$ (19)
It should be noted that when entity $i\notin\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$ the $score_{rep}^{q,i}$ is 0.
4.6 Loss Function
The model uses the cross-entropy loss function as a standard to measure loss, with the following formula:
$\mathcal{L}(score_{i},\mathcal{G}_{\mathcal{F}(i)})=\sum_{(s,r,o,t)\in\mathcal{G}_{\mathcal{F}(i)}}y_{t}\log P_{i}(o|s,r,t)$ (20)
where $P_{i}(o|s,r,t)=softmax(score_{i})$ represents the predicted probability of entities, and $y_{t}\in|\mathcal{E}|$is the label vector, where an element is 1 if the fact occurs, or 0 otherwise. Since each encoder in the model is trained separately, each one has a different loss function. When $i=rep$ , $\mathcal{F}(i)$ equals $rep$. And when $i\in\{loc,glo\}$, $\mathcal{F}(i)$ retrieves $raw$.
4.7 Final Score
The final model score combines scores from the repeating, local and global history encoders.
$\begin{split}score_{fin}^{q}=&\ \alpha\cdot score_{loc}^{q}+(1-\alpha)\cdot score_{glo}^{q}\\ &+score_{rep}^{q}\end{split}$ (21)
$\alpha\in[0,1]$ is a hyperparameter.
Model ICEWS18 ICEWS14 ICEWS05-15
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
Single-Step xERTE 29.31 21.03 33.51 46.48 40.79 32.70 45.67 57.30 46.62 37.84 52.31 63.92
RE-GCN 32.62 22.39 36.79 52.68 42.00 31.63 47.20 61.65 48.03 37.33 53.90 68.51
TITer 29.98 22.05 33.46 44.83 41.73 32.74 46.46 58.44 47.60 38.29 52.74 64.86
TiRGN 33.66 23.19 37.99 54.22 44.04 33.83 48.95 63.84 50.04 39.25 56.13 70.71
CEN 31.50 21.70 35.44 50.59 42.20 32.08 47.46 61.31 46.84 36.38 52.45 67.01
RETIA 32.43 22.23 36.48 52.94 42.76 32.28 47.77 62.75 47.26 36.64 52.90 67.76
RLGNet 34.96 24.68 39.22 55.09 46.15 36.16 51.17 65.12 50.56 40.34 56.05 70.18
Multi-Step CyGNet 26.07 16.76 29.54 44.43 34.80 25.34 39.05 53.09 38.17 27.93 43.01 57.89
RE-GCN 28.44 19.03 31.96 46.86 37.68 28.00 41.81 56.87 38.74 28.50 43.60 58.52
TiRGN 28.85 19.18 32.58 47.78 38.37 28.80 42.50 56.94 39.97 29.44 44.76 60.92
CENET 27.40 18.91 30.26 44.36 35.62 27.10 38.81 52.31 39.92 30.21 44.14 59.09
RLGNet 29.90 20.18 33.64 49.08 39.06 29.34 42.03 58.12 40.83 30.06 45.91 61.93
Table 1: Performance (in percentage) on on ICEWS18, ICESW14, and ICEWS05-15.
Model WIKI YAGO GDELT
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
Single-Step xERTE 73.60 69.05 78.03 79.73 84.19 80.09 88.02 89.78 19.45 11.92 20.84 34.18
RE-GCN 78.53 74.50 81.59 84.70 82.30 78.83 84.27 88.58 19.69 12.46 20.93 33.81
TITer 73.91 71.70 75.41 76.96 87.47 84.89 89.96 90.27 18.19 11.52 19.20 31.00
TiRGN 81.65 77.77 85.12 87.08 87.95 84.34 91.37 92.92 21.67 13.63 23.27 37.60
CEN 78.93 75.05 81.90 84.90 83.49 79.77 85.85 89.92 20.39 12.96 21.77 34.97
RETIA 78.59 74.85 81.39 84.58 81.04 77.00 83.31 88.62 20.12 12.76 21.45 34.49
RLGNet 82.43 78.86 85.65 87.17 89.69 87.05 92.15 93.00 25.09 16.95 27.42 40.87
Multi-Step CyGNet 58.44 53.03 62.24 67.46 68.60 60.97 73.58 83.16 19.11 11.90 20.31 33.12
RE-GCN 62.05 58.95 63.89 67.39 70.05 65.76 72.70 77.16 19.62 12.47 20.86 33.48
TiRGN 64.04 60.72 66.52 68.96 78.51 74.01 82.74 84.76 19.87 12.46 21.21 34.25
CENET 57.52 51.99 61.93 66.29 69.90 64.01 73.04 82.65 - - - -
RLGNet 64.34 61.03 66.71 69.51 80.17 76.52 83.57 84.96 20.81 13.34 22.32 35.38
Table 2: Performance (in percentage) on WIKI, YAGO, and GDELT.
5 Experiments
5.1 Setup
Dataset $||\mathcal{E}|$$|\mathcal{R}|$ Train Valid Test Time gap
ICE18 23,033 256 373,018 45,995 49,545 24 hours
ICE14 7,128 230 63,685 13,823 13,222 24 hours
ICE15 10,488 251 368,868 46,302 46,159 24 hours
WIKI 12,554 24 539,286 67,538 63,110 1 year
YAGO 10,623 10 161,540 19,523 20,026 1 year
GDELT 7,691 240 1,734,399 238,765 305,241 15 mins
Table 3: Statistics of the datasets.
Datasets. We use six TKG datasets to evaluate the model’s effectiveness in entity prediction task, including ICEWS14(?), ICEWS18(?), and ICEWS05-15(?) from the Integrated Crisis Early Warning System (ICEWS) and the event-driven GDELT(?) dataset. Public datasets WIKI(?) and YAGO(?) are also included. All datasets are time-partitioned into Training (80%), Validation (10%), and Test (10%). More details on datasets are in Table 3.
Evaluation Metrics. To assess TKG reasoning performance, we used Mean Reciprocal Rank (MRR) and Hits@k metrics. MRR calculates average inverse rankings for actual entities across all queries, while Hits@k denotes the proportion of real entities appearing within the top k rankings. Previous studies have indicated that traditional filtering settings are flawed(?). Thus, we report the experimental results after using the time-aware filtering settings.
Baselines. RLGNet is compared with six baseline models on single-step reasoning task, including xERTE(?), RE-GCN(?), TITer(?), TiRGN(?), CEN(?), and RETIA(?). Since some models are not designed for multi-step reasoning task, we have selected a few models to report their performance on multi-step reasoning task, including CyGNet(?), RE-GCN(?), TiRGN(?), and CENET(?).
Implementation Details. For all datasets, we set the vector dimension $d$ to 200, and the time vector dimension $d_{2}$ in the local history encoder to 48. The $top_{k}$ and $top_{k}^{all}$ are set to 20 and 200. The number of GCN layers $\omega$ is set to 1. For ICEWS18, ICEWS14, ICEWS05-15, GDELT, WIKI, and YAGO, the adjacent history length $m$ is set to 10, 10, 15, 10, 1, and 1, respectively. The hyperparameter $\alpha$ is set to 0.8 in ICEWS, 0.9 in YAGO and WIKI, and 0.1 in GDELT. At the same time, a static graph constraint similar to that of RE-GCN is also added to ICEWS. Adam is used for parameter learning, with a learning rate set to 0.001. In the local history encoder, we employ StepLR to adjust the learning rate. In this setup, the value of gamma is set to 0.8. For the YAGO and WIKI datasets, the step-size value is set to 2, while for other datasets, step-size is 10. Please note that each encoder is trained independently, which results in them not sharing weights.
5.2 Results
The reasoning results of the entity prediction task are shown in Tables 1 and 2, where RLGNet outperforms other baselines in most cases. On six benchmark datasets, especially on the GDELT and ICEWS14 datasets, the MRR scores of RLGNet have increased by 3.42% and 2.11%, respectively. The primary reason for this improvement is that RLGNet not only considers repeating, local, and global historical information, but also integrates query information, which is not simultaneously considered by other models.
In single-step reasoning performance comparison, RLGNet demonstrates stronger performance in most cases. In the TKG extrapolation task, methods that consider query information usually achieve higher H@1 results, in comparison to the methods that ignore it(?). However, these methods may perform slightly worse on other metrics like H@3 and H@10. This could be because methods that don’t consider query information may introduce answers that don’t fully match the specific query, leading to inferior performance on H@1. This is also why, in the single-step reasoning task on the ICEWS05-15 dataset, even though TiRGN surpasses RLGNet by 0.08% in H@3 and 0.53% in H@10, RLGNet still outperforms TiRGN by 1.09% in H@1.
In multi-step reasoning task, the model is unable to know the ground truth facts during the test period. Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task. Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task. This is why global historical information becomes more important in multi-step reasoning task compared to single-step reasoning. TiRGN and RE-GCN has not fully considered global historical information, while CyGNet and CENET have overlooked local historical information. In contrast, RLGNet effectively integrates repetitive, local, and global historical information based on query information. Therefore, it achieves the best performance in multi-step reasoning task.
5.3 Ablation Study
Model ICEWS18 ICWS14 ICEWS05-15 WIKI YAGO GDELT
Single Multi Single Multi Single Multi Single Multi Single Multi Single Multi
Glo 32.54 29.10 42.86 38.51 45.41 39.05 59.46 52.94 71.85 60.92 24.72 20.63
Loc 34.08 28.19 44.68 37.09 50.04 38.88 81.72 63.86 87.22 77.84 21.57 19.42
Glo + Loc 34.52 29.13 45.56 38.42 50.23 39.97 81.75 63.88 87.38 78.39 24.73 20.72
Glo + Rep 33.19 29.56 43.74 38.64 46.57 40.14 66.75 56.98 68.85 59.29 25.11 20.74
Loc + Rep 34.93 29.39 45.91 38.41 50.75 40.21 82.35 64.33 89.65 79.50 23.27 20.41
RLGNet 34.96 29.90 46.15 39.06 50.56 40.83 82.43 64.34 89.69 80.17 25.09 20.81
Table 4: The MRR (in percentage) results of the ablation studies.
$top_{k}$ ICE18 ICE14 ICE15 WIKI YAGO GDELT 5 34.49 51.94 42.77 26.69 84.05 77.67 10 40.29 58.26 47.10 34.26 90.76 85.37 20 44.63 62.90 49.82 42.35 92.63 86.96 30 46.58 64.80 50.69 47.05 92.73 87.03 100 49.76 67.59 51.75 58.41 92.73 87.04 $\infty$ 50.42 68.38 51.85 64.93 92.73 87.04
Table 5: The proportion of repeating facts when $top_{k}$ is set to different values.
To validate the impact of different modules, we conducted ablation experiments on single-step and multi-step reasoning tasks on six benchmark datasets. The results are shown in Tables 4. Our ablation studies indicated that the local history encoder (Loc) and the global history encoder (Glo) do not always improve results. Specifically, the use of Glo and Loc modules led to a decrease in performance in single-step reasoning tasks in the ICEWS05-15 and GDELT datasets. However, considering that their improvement in other tasks far outweighs the decline in these two subtasks, we thus believe Glo and Loc are beneficial. On the other hand, the Repetitive History Encoder (Rep) can improve the final performance across various tasks, suggesting that events usually exhibit a degree of repetition. As shown in Table 5, the proportion of repetitive events in the dataset well corroborates this point. Overall, while Loc and Glo may not perform ideally in some specific tasks, these three historical information encoding methods can all contribute to better prediction results.
5.4 Hyperparameter Analysis
Figure 3: The impact of $\alpha$ on MRR (in percentage) results.
Effect of $\alpha$ Values in RLGNet. When processing multi-step and single-step reasoning tasks, global and local encoders contribute differently to the answers, as shown in Figure 3. In single-step reasoning task, we find that on the ICEWS dataset, the performance is best when $\alpha$ is between 0.6 and 0.8. On the WIKI and YAGO datasets, the performance peaks when $\alpha$ is between 0.8 and 1. On the GDELT dataset, the optimum range for $\alpha$ is between 0 and 0.2. Therefore, we set $\alpha$ to 0.8 and 0.1 in ICEWS and GDELT respectively, and to 0.9 in both WIKI and YAGO. In multi-step reasoning task, in order to achieve optimal performance, the value of $\alpha$ is usually smaller than the value of $\alpha$ in single-step inference tasks. This is because the local history encoder can more easily capture the sequential pattern of facts when the baseline truth is known. It’s noteworthy that in the GDELT dataset, the contribution of global historical information always exceeds that of local historical information. This is due to the GDELT dataset containing more abstract entity concepts, making it more difficult for the model to capture the sequential patterns of facts(?).
Figure 4: The impact of $\omega$ and $top_{k}$ on MRR(in percentage) results in ICEWS14 and YAGO.
Effect of $\omega$ and $tok_{k}$ Values in RLGNet. Figure 4 shows the impact of $\omega$ and $top_{k}$ on the MRR results in the ICEWS14 and YAGO datasets. As $\omega$ increases, the performance of the model on ICEWS14 fluctuates slightly, while the performance on the YAGO dataset rapidly declines, so we set $\omega$ to 1 in all datasets. In the YAGO dataset, setting $top_{k}$ at a higher value yields better performance however, changes in $top_{k}$ do not significantly affect the performance on the ICEWS14 dataset. Therefore, we set $top_{k}$ for all datasets to 20.
6 Conclusion
In this paper, we propose RLGNet for TKG reasoning. RLGNet captures the characteristics of corresponding facts by using repeating, local, and global historical encoders respectively. In the future, finding a superior strategy to balance the contributions of global and local historical information is a worthwhile research problem. Experiments on six benchmark datasets prove that RLGNet outperforms existing models in most extrapolation tasks.
7 Limitations
In this paper, we have proposed a TKG prediction model based on three encoders. However, our approach has certain limitations:
  • 1.
    Due to the constraints of computational resources, the three encoders need to train independently. This method prevents them from sharing weights among themselves, potentially affecting the predictive performance of the model.
  • 2.
    In multi-step and single-step reasoning tasks, the contributions of global and local historical information should indeed be different. We have adopted a weighted average strategy to alleviate this issue to some extent, but choosing a specific hyperparameter $\alpha$ cannot achieve optimal performance on both multi-step and single-step reasoning tasks at the same time.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No.61866008).

### v2

@@
 Abstract
-Temporal Knowledge Graph (TKG) reasoning is based on historical information to predict the future.
-Therefore, parsing and mining historical information is key to predicting the future.
-Most existing methods fail to concurrently address and comprehend historical information from both global and local perspectives.
-Neglecting the global view might result in overlooking macroscopic trends and patterns, while ignoring the local view can lead to missing critical detailed information.
-Additionally, some methods do not focus on learning from high-frequency repeating events, which means they may not fully grasp frequently occurring historical events.
-To this end, we propose the Repetitive-Local-Global History Network(RLGNet).
-We utilize a global history encoder to capture the overarching nature of historical information.
-Subsequently, the local history encoder provides information related to the query timestamp.
-Finally, we employ the repeating history encoder to identify and learn from frequently occurring historical events.
+Temporal Knowledge Graph (TKG) reasoning involves predicting future events based on historical information.
+However, due to the unpredictability of future events, this task is highly challenging.
+To address this issue, we propose a multi-scale hybrid architecture model based on ensemble learning, called RLGNet (Repeating-Local-Global History Network).
+Inspired by the application of multi-scale information in other fields, we introduce the concept of multi-scale information into TKG reasoning.
+Specifically, RLGNet captures and integrates different levels of historical information by combining modules that process information at various scales.
+The model comprises three modules: the Repeating History Module focuses on identifying repetitive patterns and trends in historical data, the Local History Module captures short-term changes and details, and the Global History Module provides a macro perspective on long-term changes.
+Additionally, to address the limitations of previous single-architecture models in generalizing across single-step and multi-step reasoning tasks, we adopted architectures based on Recurrent Neural Networks (RNN) and Multi-Layer Perceptrons (MLP) for the Local and Global History Modules, respectively.
+This hybrid architecture design enables the model to complement both multi-step and single-step reasoning capabilities.
+Finally, to address the issue of noise in TKGs, we adopt an ensemble learning strategy, combining the predictions of the three modules to reduce the impact of noise on the final prediction results.
 In the evaluation on six benchmark datasets, our approach generally outperforms existing TKG reasoning models in multi-step and single-step reasoning tasks.
+Figure 1: The figure illustrates the repeating, local, and global perspectives. The red timestamp indicates that an fact identical to the queried fact occurred at that specific moment in time.
 1 Introduction
-TKG is a structured yet highly complex knowledge system.
-In TKG, each fact is represented by a quadruple, including subject, relation, object, and timestamp, thereby providing detailed and accurate temporal information.
-TKG allow us to understand the relationships between entities and capture the changes in these relationships via timestamps.
-TKG extrapolation aims to predict new facts at a future timestep based on the ground truth facts within the known timeframe.
-Extrapolation is further classified into two types: (1) Single-step reasoning predicts future events using known ground truth facts from past timestamps within the test period.
-(2) Multi-step reasoning forecasts without relying on ground truth facts from the test period.
-By analyzing past events and identifying factors that influence event progression, we can infer the probabilities of emergent future facts.
-However, since future events contain unknown factors, TKG extrapolation is a challenging problem(?).
-Figure 1: A figure to explain repeating, local, and global facts.
-”Frequency” represents the number of occurrences of that type of event, ”Interval” represents the time difference from the last occurrence of that type of event to the time of query.
-The red arrows represent the facts most relevant to the query.
-Predicting the future is a task that requires a deep analysis of historical data, the key to which is to discern possible patterns and trends from it.
-In previous work, it’s generally believed that the facts closer to the query time are more important.
-However, this method may cause us to overlook some facts that have the potential to impact future trends, because these facts are not necessarily always close to the query time.
-As shown in Figure 1, (United States, Negotiations, ?) does not appear in the local history.
-According to historic recurrence(?), historical facts could have repeated or even cyclical patterns.
-As illustrated in the same Figure 1, (United States, Negotiations, North Korea) has occurred 95 times.
-Therefore, we need to consider the repetitiveness, locality, and globality of historical information.
-Moreover, some work has shown that not every fact in the TKG is crucially important for prediction.
-To improve the accuracy of prediction, we need to selectively pay attention to the facts that are most relevant to the query, as the fact indicated by the red arrow in Figure 1.
-Therefore, finding a method that is based on the query and simultaneously considers repeating, local and global historical information is of paramount importance for extrapolation in TKG.
-To address this challenge, we design and propose a model called Repeating-Local-Global History Network (RLGNet).
-In this model, we design repeating, local, and global history encoders respectively based on the query-oriented method.
-In the local history encoder, the model views the TKG as a Knowledge Graph (KG) sequence and models the facts in adjacent timestamps to obtain local historical information.
-In the global history encoder, the model aggregates all facts using an attention mechanism to gather global historical information.
-Finally, building on the foundation of local and global historical encoders, we use a repeating history encoder to further increase the scores of repetitive events.
-Overall, this paper makes the following contributions:
+A Temporal Knowledge Graph (TKG) is a structured yet highly complex knowledge system.
+In TKG, each fact is represented by a quadruple consisting of subject, relation, object, and timestamp.
+This paper primarily discusses the problem of extrapolation in TKG.
+TKG allows us to understand the relationships between entities and capture changes in these relationships over time.
+The goal of TKG extrapolation is to predict new facts at future time steps based on known facts.
+Extrapolation can be further divided into single-step and multi-step types.
+Single-step reasoning involves predicting facts that will occur at the current time step (time $t$) based on known past facts.
+Multi-step reasoning requires sequentially inferring facts at multiple future time steps, such as $t$, $t+1$, $t+2$, and so on.
+However, the inclusion of unknown factors in future events poses significant challenges for TKG extrapolation (?).
+Recent studies indicate that some facts in TKG exhibit traceable patterns, hence simulating the evolution of these facts can aid in predicting them.
+For instance, models based on RNN architecture are widely used (?).
+Additionally, there are models based on MLP (?), where time is often vectorized and input as a feature to capture the temporal characteristics of facts.
+To achieve more accurate predictions, some models cleverly design methods to fuse local and global historical information, allowing the model to have both global and local perspectives.
+The global perspective helps the model better generalize the entire dataset’s characteristics, while the local perspective assists in extracting recent fact information.
+Furthermore, some studies identify closely related entities as candidate answers, using them to constrain the prediction range and improve accuracy.
+Beyond these designs, we further observe that a portion of facts in TKG are closely related to the past, as evidenced by the significant proportion of facts that repeat multiple times in TKG.
+Thus, when predicting the future, it is reasonable to assume that these facts will reoccur.
+Conversely, some facts in TKG only occur at specific times, contributing to substantial noise in TKG datasets.
+This noise causes models to fit these time-specific facts during training, impairing their ability to predict future events.
+In TKG multi-step reasoning tasks, the absence of facts over a period can cause RNN-based models to evolve incorrect facts.
+Over time, the influence of past facts diminishes, and MLP-based models tend to assign similar weights to facts at different times, affecting their performance in single-step reasoning.
+Just as RNN and MLP models each have their strengths in single-step and multi-step reasoning tasks, they also have advantages in capturing historical information at different scales.
+MLP-based models are well-suited for capturing patterns and relationships across the entire dataset because they process each input independently, without temporal bias, allowing them to generalize long-term trends and patterns.
+On the other hand, RNN-based models are ideal for capturing local historical information, as they excel at handling short sequences of data and can effectively capture short-term changes in the most recent time steps.
+Based on these observations, we designed a model called the Repeating-Local-Global History Network (RLGNet).
+Specifically, RLGNet draws inspiration from ensemble learning by designing three relatively independent modules: Global History Module, Local History Module, and Repeating History Module.
+As shown in Figure 1, these three modules capture information at different scales.
+During training, RLGNet first learns the Local and Global History Modules in parallel, while the Repeating History Module learns serially on the results of the Global and Local History Modules to further reinforce the scores of repeating facts.
+This ensemble learning strategy can reduce the sensitivity of individual models to noisy data in TKG, better capture the true patterns in the data, and reduce the risk of overfitting.
+Due to the advantages of RNN-based and MLP-based models in single-step and multi-step reasoning tasks, as well as their effectiveness in handling historical information at different scales, RLGNet adopts an RNN-based structure for the Local History Module and an MLP-based structure for the Global History Module.
+This hybrid design fully leverages the complementary strengths of both approaches.
+Finally, to further constrain the answer range, we integrate candidate entities’ information into the three learning processes.
+Our multi-scale hybrid architecture design enables RLGNet to achieve state-of-the-art performance in most cases for both multi-step and single-step reasoning tasks.
+Our contributions are summarized as follows:
   1. 1.
-    We designed three different levels of encoders to deal with the repetitive, local, and global historical information respectively.
+    Unlike previous work, we designed an ensemble learning-based model that processes different historical information from a multi-scale perspective.
+Specifically, we use Repeating, Local, and Global History Modules to capture their corresponding historical information, and these modules collaborate to enhance the overall performance of the model.
   2. 2.
-    We explored the contributions of repeating, global and local historical information to predictive performance in two different extrapolation settings.
+    We use RNN and MLP architectures in the Local History Module and Global History Module, respectively, and integrate their scores to achieve complementary capabilities between the two modules.
+This hybrid architecture model addresses the limitation of previous works, where a single architecture could only achieve good results in either multi-step or single-step tasks, improving the model’s performance in multi-task scenarios.
   3. 3.
-    We conducted extensive experiments on six public TKG datasets, proving the model’s effectiveness in entity prediction.
+    We proposed a multi-scale hybrid architecture model based on an ensemble learning strategy, which demonstrated impressive performance across six public TKG datasets.
 2 Related Work
-Static KG Reasoning: Recent years have seen great interest and research in static Knowledge Graph (KG) reasoning models.(??) These models include distance-based models, such as TransE(?) and TransH(?), which determine the likelihood of facts by measuring the distances between entities.
+Static KG Reasoning.
+Recent years have seen great interest and research in static Knowledge Graph (KG) reasoning models.(??) These models include distance-based models, such as TransE(?) and TransH(?), which determine the likelihood of facts by measuring the distances between entities.
 Another category is semantic-matching-based models, like DistMult(?) and RESCAL(?).
@@
 However, these models focus on static KG, and their predictive capabilities for future events are limited.
-Temporal KG Reasoning: TKG reasoning has two settings: extrapolation and interpolation.
+Temporal KG Reasoning.
+TKG reasoning has two settings: extrapolation and interpolation.
 The interpolation setting aims at predicting missing historical facts rather than future events.
@@
 3 Preliminaries
-TKG represent events temporally as snapshots.
-Let $\mathcal{G}_{raw}$ be the sequence of snapshot graphs $\mathcal{G}_{raw}=\{\mathcal{G}_{1},\mathcal{G}_{2},\dots,\mathcal{G}_{T}\}$.
+TKGs represent events temporally as snapshots.
+Let $\mathcal{G}_{raw}$ be the sequence of snapshot graphs $\mathcal{G}_{raw}=\{\mathcal{G}_{raw}^{1},\dots,\mathcal{G}_{raw}^{|\mathcal{T}|}\}$.
 Each fact in the snapshot is a quadruple $(s,r,o,t)$, where $s,o\in\mathcal{E},r\in\mathcal{R},t\in\mathcal{T}$.
 It represents a relationship $r$ between subject entity $s$ and object entity $o$ at time $t$.
-For each quadruple $(s,r,o,t)$,an inverse relation quadruple$(o,r^{-1},s,t)$ is often added to the dataset.
-The entity prediction task aims to predict the missing entity in a query $q=(s_{q},r_{q},?,t_{q})$.
-We denote all entity embeddings in dimension $d$ with $H\in\mathbb{R}^{|\mathcal{E}|\times d}$, and relation embeddings similarly with $R\in\mathbb{R}^{|\mathcal{R}|\times d}$.
-In each encoder, the initial weights of H and R are randomly generated.
-The specific entity and relation are represented as $h\in\mathbb{R}^{d}$and $r\in\mathbb{R}^{d},$ respectively.
-Given the query tuple $q=(s_{q},r_{q},?,t_{q})$, we define the candidate entity set at time $t$ as $\mathcal{C}_{q}^{t}=\{o|(s_{q},o_{q},o,t)\in\mathcal{G}_{t}\}$.
-The candidate entity set over a time range $t_{1}$ to $t_{2}$ is defined as $\mathcal{C}_{q}^{t_{1}:t_{2}}$, a union of all candidate sets over the range.
+$\mathcal{E}$, $\mathcal{R}$, and $\mathcal{T}$ represent the sets of entities, relations, and timestamps, respectively.
+In this paper, our reasoning task is to predict missing entities, specifically the subject entity $s$ and the object entity $o$.
+To simplify the problem, we convert the task of predicting the object entity $o$ into predicting the subject entity $s$, represented as $(o,r^{-1},?,t)$.
+Therefore, for each quadruple $(s,r,o,t)$, we add a reverse relation quadruple $(o,r^{-1},s,t)$ to the dataset.
+Given the query tuple $q=(s_{q},r_{q},?,t_{q})$, we define the candidate entity set at time $t$ as $\mathcal{C}_{q}^{t}=\{o|(s_{q},r_{q},o,t)\in\mathcal{G}^{t}_{raw}\}$.
+The candidate entity set over a time range $t_{1}$ to $t_{2}$ is defined as $\mathcal{C}_{q}^{t_{1}:t_{2}}$, a union of all candidate sets over the range:
 $\mathcal{C}_{q}^{t_{1}:t_{2}}=\bigcup_{i=t_{1}}^{t_{2}}\mathcal{C}_{q}^{i}$ (1)
-Since not all candidate entities greatly affect the prediction, we only keep the top $k$ most frequent entities, denoted as $C_{q,top_{k}}^{1:t_{q}-1}$.
-Candidates in the set usually represent repeating facts, we construct $\mathcal{G}_{rep}$, containing all tuples in the period where o is a top $k$ candidate before $t-1$.
-$\mathcal{G}_{rep}$ can be formalized as:
-$\mathcal{G}_{rep}=\{q=(s_{q},r_{q},o_{q},t_{q})\in\mathcal{G}_{raw}|o\in\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}\}$ (2)
-$\mathcal{G}_{rep}$ essentially extracts prominent, repetitive events from historical data.
-Figure 2: The overall architecture of RLGNet.
-The upper half is the local history encoder which learns local historical information through KG sequences of adjacent timestamps.
-The lower half comprises the repeating and global history encoders, learning repetitive and global historical information respectively by statistically querying candidate entities.
-The content within the dashed box represents the entity most relevant to the query.
-The left side sums the scores of the three parts to get the final prediction score.
+Since not all candidate entities greatly affect the prediction, we only keep the top $top_{k}$ most frequent entities, denoted as $\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$.
+To represent repeating facts, we need to select those facts that have occurred more than once in history and define repeating facts that occurred at time $t$ as $\mathcal{G}_{rep}^{t}$.
+$\mathcal{G}_{rep}^{t}$ can be formalized as:
+$\mathcal{G}_{rep}^{t}=\{q^{\prime}|q^{\prime}=(s,r,o,t)\in\mathcal{G}^{t}_{raw},o\in\mathcal{C}_{q^{\prime},top_{k}}^{1:t-1}\}$ (2)
+Similar to $\mathcal{G}_{raw}$, we obtain $\mathcal{G}_{rep}=\{\mathcal{G}_{rep}^{1},\dots,\mathcal{G}_{rep}^{|\mathcal{T}|}\}$.
+Figure 2: The upper half represents the Local History Module, which learns local historical information through KG sequences of adjacent timestamps.
+The lower half comprises the Repeating and Global History Modules, learning repetitive and global historical information respectively by statistically querying candidate entities.
+The entities within the dashed box are candidate entities.
+The left side weights and sums the scores of the three modules to obtain the final prediction score.
 4 Model Overview
-The overall framework of the model is illustrated in the Figure 2.
-This model can be divided into three core submodules: the local history encoder, the global history encoder and the repeating history encoder.
-The local history encoder captures local facts by exploring the structural features and historical dependencies the sequence of KG at adjacent timestamps.
-The global history encoder captures global facts by examining the relevant facts across all previous timestamps and capturing entities and relationships that may not be evident at adjacent timestamps.
-Furthermore, these two encoders use a scoring decoder to evaluate the facts.
-The primary task of the repeating history encoder is to enhance its ability to predict repetitive facts by encoding frequently occurring historical events.
-All encoders utilize time vector to encode time and frequency.
-By utilizing these three encoders, the model is capable of making accurate predictions in response to queries.
-4.1 Scoring Decoder
-Inspired by the work of ConvE, we designed this scoring decoder and used it in the next two decoders.
-First, the model reshapes a one-dimensional input into a two-dimensional matrix.
-Then, it processes this matrix with a CNN.
-After that, it flattens the resulting data and outputs the embeddings for the target entity through a fully connected layer.
-$\varphi(H,V_{1}\dots V_{k})=\sigma(ConvE(V_{1},\dots,V_{k}))\cdot H$ (3)
-$\sigma$ represents the ReLu function. $V_{i}(1\leq i\leq k)$are input vectors, $\varphi(H,V_{1},\dots,V_{k})\in\mathbb{R}^{|\mathcal{E}|}$ represents predicted entity scores.
-4.2 Time Vector
-In terms of time, the module employs two types of temporal encoding vectors, periodic(?) and aperiodic, to represent time.
-$\textbf{v}_{1}(t)&=\cos(W_{1}^{t}t+b_{1})\\ \textbf{v}_{2}(t)&=\tanh(W_{2}^{t}t+b_{2})\end{aligned}\right.$ (4)
-Then we concatenate the two vectors to obtain the final vector representation.
-$\textbf{V}(t)=[\textbf{v}_{1}(t);\textbf{v}_{2}(t)]$ (5)
-$W_{1}^{t},W_{2}^{t},b_{1},b_{2}$ are learnable parameters and the vectors $\textbf{v}_{1}(t)$ and $\textbf{v}_{2}(t)$ have the same dimensionality, [;] denotes the vector concatenation operation, $\textbf{V}(t)$ is the time vector.
-4.3 Local History Encoder
+The overall framework of the model is illustrated in Figure 2.
+This model can be divided into three core submodules: the Local History Module, the Global History Module, and the Repeating History Module.
+The Local and Global History Modules are used to capture historical information at local and global scales, respectively, with the Local History Module based on an RNN architecture and the Global History Module based on an MLP architecture.
+The Repeating History Module is used to further enhance the scores of repeating facts.
+To handle the frequency of time and entity occurrences, we use a Numerical Embedding function to convert them into embeddings.
+To further constrain the answer range, we apply the candidate entity design to all three modules.
+The model is designed based on the concept of ensemble learning, so the repeating, global, and local modules are trained independently, and their prediction scores are combined through weighted summation to produce the final result.
+4.1 Numerical Embedding
+The Numerical Embedding module is used to convert time and entity occurrence frequency into embeddings.
+The cosine function can capture periodic patterns in time and frequency, while the tanh function can provide smoother transitions for phenomena with rapidly changing frequencies or sudden changes in time.
+$\textbf{v}_{1}(x)&=\cos(W_{1}^{N}x+b_{1}^{N})\\ \textbf{v}_{2}(x)&=\tanh(W_{2}^{N}x+b_{2}^{N})\end{aligned}\right.$ (3)
+Then we concatenate the two embeddings to obtain the final embedding.
+$\textbf{V}(t)=[\textbf{v}_{1}(x);\textbf{v}_{2}(x)]$ (4)
+Here, $W_{1}^{N},W_{2}^{N},b_{1}^{N},b_{2}^{N}$ are learnable parameters and the embeddings $\textbf{v}_{1}(x)$ and $\textbf{v}_{2}(x)$ have the same dimensionality.
+The notation [;] denotes the embedding concatenation operation, and $\textbf{V}(x)$ represents the Numerical Embedding.
+Combining the embeddings generated by these two modules with different characteristics allows for the integration of both periodic and nonlinear information, providing a more comprehensive representation.
+4.2 Local History Module
 This module captures local facts by focusing on adjacent history.
-For each query $q=(s_{q},r_{q},?,t_{q})$, the module pays attention to the $m$ timestamp subgraph $\{\mathcal{G}_{t_{q}-m},\dots,\mathcal{G}_{t_{q}-1}\}$ related to this query to obtain the structural features of the subgraph.
-For this purpose, we employ GCN to aggregate individual subgraphs, use Gated Recurrent Units (GRU)(?) to learn the features of subgraph evolution, and adopt an attention mechanism to integrate information from candidate entities.
-Firstly, we use GCN with $\omega$ layers to get the entity representations at the current moment.
-$h_{o}^{t,l+1}=\sum\limits_{(s,r,o)\in\mathcal{G}_{t}}\frac{1}{|N_{o}^{t}|}W_{1}^{l}\phi(h_{i}^{t,l},r_{t})+W_{2}^{l}h_{o}^{l}$ (6)
-$\phi(h_{i}^{t,l},r_{t})=[h_{i}^{t,l};r_{t};h_{i}^{t,l}+r_{t};h_{i}^{t,l}\cdot r_{t}]$ (7)
-The symbol [$\cdot$] represents the Hadamard product.
-$N_{o}^{t}$ represents neighbors of node $o$ at time $t$.
-The embedded representation of node $o$ at the $l$-th layer is denoted as $h_{o}^{t,l}\in\mathbb{R}^{d}$.
-The aggregation and self-loop parameters of the $l$-th layer are shown as $W_{1}^{l}$ and $W_{2}^{l}$.
-The edge relation type is presented as $r_{t}$.
-Specifically, when $\omega=0$, i.e., $h_{o}^{t,0}=\sum_{i\in N_{o}^{t}}\frac{1}{|N_{o}^{t}|}h_{i}^{t}$.
-Entity representations are combined with their temporal representations and then predicted for the next moment with GRU:
-$H_{t+1}=GRU([H_{t};\textbf{V}_{1}(t_{q}-t)],H^{{}^{\prime}})$ (8)
-$\textbf{V}_{1}(t_{q}-t)\in\mathbb{R}^{d_{2}}$ is the time vector.
-$H^{{}^{\prime}}\in\mathbb{R}^{|\mathcal{E}|\times d}$ is the final GCN layer output.
+For each query $q=(s_{q},r_{q},?,t_{q})$, the module pays attention to the $m$ timestamp subgraph $\{\mathcal{G}^{t_{q}-m}_{raw},\dots,\mathcal{G}^{t_{q}-1}_{raw}\}$ related to this query to obtain the structural features of the subgraph.
+Here, $m$ is a hyperparameter.
+To achieve this, we employ a GCN to aggregate individual subgraphs, use Gated Recurrent Units (GRU) (?) to learn the features of subgraph evolution, and adopt an attention mechanism to integrate information from candidate entities.
+Firstly, we use a GCN with $\omega$ layers to get the entity embeddings at the current moment:
+$h_{o}^{t,l+1}=\sum\limits_{(s,r,o,t)\in\mathcal{G}^{t}_{raw}}\frac{1}{|N_{o}^{t}|}W_{1}^{l}\phi(h_{s}^{t,l},r^{L})+W_{2}^{l}h_{o}^{l}$ (5)
+$\phi(h_{s}^{t,l},r^{L})=[h_{s}^{t,l};r^{L};h_{s}^{t,l}+r^{L};h_{s}^{t,l}\cdot r^{L}]$ (6)
+Here, the symbol [$\cdot$] represents the Hadamard product.
+$N_{o}^{t}$ represents the neighbors of node $o$ at time $t$.
+The embedded of node $o$ and $s$ at the $l$-th layer is denoted as $h_{o}^{t,l}$ and $h_{s}^{t,l}$.
+The aggregation and self-loop parameters of the $l$-th layer are denoted as $W_{1}^{l}$ and $W_{2}^{l}$.
+$r^{L}$ is the initial embedding of the relation in the Local History Module.
+When $\omega=0$,$h_{o}^{t,0}=\sum_{i\in N_{o}^{t}}\frac{1}{|N_{o}^{t}|}h_{i}^{t}$.
+If $t=t_{q}-m$ then $h^{t_{q}-m}_{o}$ represents the initial embedding of the entity $o$ in the Local History Module.
+Entity embeddings are combined with their temporal embeddings and then predicted for the next moment with a GRU:
+$H_{t+1}=GRU([H_{t};\textbf{V}^{L}(t_{q}-t)],H^{{}^{\prime}}_{t})$ (7)
+Here, $\textbf{V}^{L}$ is Numerical Embeddings.
+$H^{{}^{\prime}}_{t}$ is the output of the last layer of the GCN for all entities at time $t$, and $H_{t}$ is the embedding representation of all entities at time $t$.
 The attention mechanism aggregates candidate entities over a history of length $m$.
 Node aggregation for query $q$ is:
-$C_{q}^{t}=\sum\limits_{i\in C_{q}^{t}}\frac{1}{|C_{q}^{t}|}h_{i}^{{}^{\prime}}$ (9)
-$C_{q}^{t}$ represents query $q$’s candidate entity representation at time $t$. Next, the module calculates attention weights $a_{q}^{t}$ for the query at time $t$:
-$a_{q}^{t}=W_{2}\sigma(W_{1}[h_{q}^{{}^{\prime}};r_{q};\textbf{V}_{1}(t_{q}-t);C_{q}^{t}])$ (10)
-$W_{1}$, and $W_{2}$ are learnable parameters. After calculating attention weights, these are used to aggregate candidate entities for query q:
-$C_{q}^{{}^{\prime}}=\sum_{i=1}^{m}\frac{\exp{(a_{q}^{i})}\cdot C_{q}^{i}}{\sum_{j=1}^{m}\exp{(a_{q}^{j})}}$ (11)
-The candidate entity representation $C_{q}^{{}^{\prime}}$ for query $q$. The module then calculates the entity score for query $q$’s local history encoder:
-$score_{loc}^{q}=\varphi_{loc}(H^{{}^{\prime}},h_{q}^{{}^{\prime}},r_{q},C_{q}^{{}^{\prime}})$ (12)
-Eq (12) can be viewed as an instance of Eq (3).
-4.4 Global History Encoder
-The global history encoder is designed to consider and extract information from a global perspective in order to capture global facts.
-For each query, we count the frequencies and the recent appearances of candidate entities.
+$C_{q}^{t}=\sum\limits_{i\in\mathcal{C}_{q}^{t}}\frac{1}{|\mathcal{C}_{q}^{t}|}h_{i}^{t}$ (8)
+Here, $C_{q}^{t}$ represents query $q$’s candidate entity embedding at time $t$.
+$h_{i}^{t}$ represents the embedding of candidate entity $i$ in $H_{t}$.
+Next, the module uses a 2-layer MLP to calculate the attention weights $a_{q}^{t}$ for the query at time $t$:
+$a_{q}^{t}=MLP([h_{q}^{t};r_{q}^{L};\textbf{V}^{L}(t_{q}-t);C_{q}^{t}])$ (9)
+Here, $r_{q}^{L}$ is the initial embedding of the query relation in the Local History Module.
+$h_{q}^{t}$ represents the embedding of the query entity in $H_{t}$.
+After calculating attention weights, these are used to aggregate candidate entities for query q:
+$C_{q}^{{}^{\prime}}=\sum_{i=t_{q}-m}^{t_{q}-1}\frac{\exp{(a_{q}^{i})}\cdot C_{q}^{i}}{\sum_{j=t_{q}-m}^{t_{q}}\exp{(a_{q}^{j})}}$ (10)
+This results in the candidate entity embedding $C_{q}^{{}^{\prime}}$ for query $q$.
+The Local History Module uses ConvTransE as the scoring function.
+To better adapt to the task, we additionally concatenate query-related information embeddings to the original entity and relation embeddings in ConvTransE.
+These concatenated embeddings are then processed through a one-dimensional convolution layer followed by a fully connected layer to obtain the scores:
+$score_{loc}^{q}=H_{t_{q}}\cdot\phi_{L}(h_{q}^{t_{q}},r_{q}^{L},C_{q}^{{}^{\prime}})$ (11)
+Here, $\phi_{L}$ represents the score decoder based on ConvTransE in the Local History Module.
+4.3 Global History Module
+The Global History Module is designed to consider and extract information from a global perspective in order to capture global facts.
+For each query, we calculate the occurrence frequency of candidate entities and the last time they appeared.
 The module then uses an attention mechanism to integrate this information about the candidate entities.
-First, the attention scores for the time-related candidate entities of the current query are calculated:
-$\mathcal{A}_{1}(q,i)=\sigma(W_{3}[h_{q};r_{q}])^{T}\sigma(W_{4}[\textbf{V}_{2}(t_{q}-t_{q}^{i});h_{i}])$ (13)
-The attention scores for frequency-related candidates is similarly:
-$\mathcal{A}_{2}(q,i)=\sigma(W_{5}[h_{q};r_{q}])^{T}\sigma(W_{6}[\textbf{V}_{3}(cnt^{i}_{q});h_{i}])$ (14)
-Considering the similarities between the characteristics of frequency and time, we employ a time vector to represent frequency.
-$W_{4}\in\mathbb{R}^{d\times 2d},W_{3}\in\mathbb{R}^{d\times 2d},W_{5}\in\mathbb{R}^{d\times 2d}$, and $W_{6}\in\mathbb{R}^{d\times 2d}$ are learnable parameters.
-$t_{q}^{i}$ and $cnt_{q}^{i}$ represent the frequency and last occurrence of candidate entity $i$ for query q.
-Additionally, taking computational resources into consideration, we only selected $top_{k}^{all}$ candidate entities.
-Following, we normalize their respective scores and multiply by the candidate entity representations to obtain representations $C_{q}^{gap}$ and $C_{q}^{cnt}$ for time-relevant and frequency-relevant candidates, respectively.
-$C_{q}^{gap}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{1}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{1}(q,j))}$ (15)
-$C_{q}^{cnt}=\sum_{i\in\mathcal{C}_{q}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{2}(q,i))\cdot h_{i}}{\sum\limits_{j\in\mathcal{C}_{q}^{1:t_{q}-1}}\exp(\mathcal{A}_{2}(q,j))}$ (16)
-The module then calculates the entity score for the global history encoder:
-$score_{glo}^{q}=\varphi_{glo}(H,h_{q},r_{q},\textbf{V}_{4}(t_{q});C_{q}^{gap};C_{q}^{cnt})$ (17)
-$h_{q}$ and $r_{q}$ represent the query entity and relation, Eq (17) can be viewed as an instance of Eq (3).
-4.5 Repeating History Encoder
-The Repeating History Encoder is designed to improve prediction by increasing the weight of historical events.
-We first filter out and retain all the repetitive facts.
-Then, only the candidate entities of facts that rank in the top($top_{k}$) in terms of frequency are retained.
-In this module, the frequency of the candidate entities is also represented using a time vector.
-The repeating history encoder’s score function is a multilayer perceptron (MLP):
-$\textbf{MLP}(x)=W_{7}\cdot\sigma(W_{8}\cdot\sigma(W_{9}\cdot x))$ (18)
-$W_{7}\in\mathbb{R}^{1\times d}$, $W_{8}\in\mathbb{R}^{d\times 3d}$, and $W_{9}\in\mathbb{R}^{3d\times 5d}$ are learnable parameters. The extra score for candidates is calculated as follows:
-$score_{rep}^{q,i}=\textbf{MLP}([h_{q};r_{q};\textbf{V}_{5}(t_{q});h_{i};\textbf{V}_{6}(cnt_{q}^{i})])$ (19)
-It should be noted that when entity $i\notin\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$ the $score_{rep}^{q,i}$ is 0.
+$\mathcal{A}_{1}(q,i)=(W_{1}^{G}[h_{q}^{G};r_{q}^{G}])^{T}(W_{2}^{G}[\textbf{V}_{1}^{G}(t_{q}-t_{q}^{i});h_{i}^{G}])$ (12)
+The attention scores for frequency-related candidates are calculated similarly:
+$\mathcal{A}_{2}(q,i)=(W_{3}^{G}[h_{q}^{G};r_{q}^{G}])^{T}(W_{4}^{G}[\textbf{V}_{2}^{G}(cnt^{i}_{q});h_{i}^{G}])$ (13)
+where $h_{q}^{G}$, $r_{q}^{G}$, and $h_{i}^{G}$ refer to the initial embeddings of the query entity, relation, and candidate entity $i$ in the Global History Module.
+$\textbf{V}_{1}^{G}$ and $\textbf{V}_{2}^{G}$ are Numerical Embeddings.
+$W_{1}^{G}$, $W_{2}^{G}$, $W_{3}^{G}$, and $W_{4}^{G}$ are learnable parameters.
+$t_{q}^{i}$ and $cnt_{q}^{i}$ represent the last occurrence and frequency of candidate entity $i$ for query $q$.
+Additionally, taking computational resources into consideration, we only selected the $top_{k}^{all}$ candidate entities.
+Following this, we normalize their respective scores and multiply by the candidate entity embeddings to obtain the embedding $C_{q}^{gap}$ for time-relevant:
+$C_{q}^{gap}=\sum_{i\in\mathcal{C}_{q,top_{k}^{all}}^{1:t_{q}-1}}\frac{\exp(\mathcal{A}_{1}(q,i))\cdot h_{i}^{G}}{\sum\limits_{j\in\mathcal{C}_{q,top_{k}^{all}}^{1:t_{q}-1}}\exp(\mathcal{A}_{1}(q,j))}$ (14)
+We normalize $\mathcal{A}_{2}$ and multiply it by the entity embeddings, then sum the results to obtain the frequency-related candidate embedding $C_{q}^{cnt}$, just as we do with $C_{q}^{gap}$.
+Similar to Equation 11, the score for the Global History Module is:
+$score_{Glo}^{q}=H_{G}\cdot\phi_{G}(h_{q}^{G},r_{q}^{G},\textbf{V}_{3}^{G}(t_{q}),C_{q}^{gap},C_{q}^{cnt})$ (15)
+where $H_{G}$ represents the initial embeddings of all entities in the Global History Module.
+$\textbf{V}_{3}^{G}$ is Numerical Embeddings.
+$\phi_{G}$ represents the score decoder based on ConvTransE in the Global History Module.
+4.4 Repeating History Module
+The Repeating History Module is designed to improve prediction by increasing the weight of historical events.
+We first filter and retain the repeated facts, and refer to the collection of these repeated facts as $\mathcal{G}_{rep}$.
+Then, only the candidate entities of facts that rank in the top $top_{k}$ in terms of frequency are retained.
+The Repeating History Module’s score function is a three-layer MLP. The candidate entities’ scores are calculated using the MLP as follows:
+$score_{Rep}^{q,i}=MLP([h_{q}^{R};r_{q}^{R};\textbf{V}_{1}^{R}(t_{q});h_{i}^{R};\textbf{V}_{2}^{R}(cnt_{q}^{i})])$ (16)
+where $h_{q}^{R}$,$h_{i}^{R}$ and $r_{q}^{R}$ represent the initial embeddings of the query entity, candidate entity, and relation in the Repeating History Module, respectively.
+$\textbf{V}_{1}^{R}$ and $\textbf{V}_{2}^{R}$ are Numerical Embeddings.
+It should be noted that when candidate entity $i\notin\mathcal{C}_{q,top_{k}}^{1:t_{q}-1}$, the $score_{Rep}^{q,i}$ is 0.
 4.5 Loss Function
-The model uses the cross-entropy loss function as a standard to measure loss, with the following formula:
-$\mathcal{L}(score_{i},\mathcal{G}_{\mathcal{F}(i)})=\sum_{(s,r,o,t)\in\mathcal{G}_{\mathcal{F}(i)}}y_{t}\log P_{i}(o|s,r,t)$ (20)
-where $P_{i}(o|s,r,t)=softmax(score_{i})$ represents the predicted probability of entities, and $y_{t}\in|\mathcal{E}|$is the label vector, where an element is 1 if the fact occurs, or 0 otherwise.
-Since each encoder in the model is trained separately, each one has a different loss function.
-When $i=rep$ , $\mathcal{F}(i)$ equals $rep$.
-And when $i\in\{loc,glo\}$, $\mathcal{F}(i)$ retrieves $raw$.
+Since the model employs an ensemble learning strategy, each module is trained independently and uses the cross-entropy loss function to calculate the loss value:
+$\mathcal{L}(score,\mathcal{G})=\sum_{\mathcal{G}^{t}\in\mathcal{G}}\sum_{(s,r,o,t)\in\mathcal{G}^{t}}y_{t}\log P(o|s,r,t)$ (17)
+The score represents the predicted scores of each module, and $\mathcal{G}$ represents the sequence of snapshot graphs.
+Here, $P(o|s,r,t)=\text{softmax}(score)$ represents the predicted probability of entities, and $y_{t}\in|\mathcal{E}|$ is the label vector, where an element is 1 if the fact occurs, or 0 otherwise.
+Therefore, the loss functions for the three modules are $\mathcal{L}(score_{Loc},\mathcal{G}_{raw})$, $\mathcal{L}(score_{Glo},\mathcal{G}_{raw})$, and $\mathcal{L}(score_{Rep},\mathcal{G}_{rep})$, respectively.
 4.6 Final Score
-The final model score combines scores from the repeating, local and global history encoders.
+The final model score combines scores from the Repeating, Local, and Global History Modules:
 $\begin{split}score_{Fin}^{q}=&\ \alpha\cdot score_{Loc}^{q}+(1-\alpha)\cdot score_{Glo}^{q}\\ &+score_{Rep}^{q}\end{split}$ (18)
-$\alpha\in[0,1]$ is a hyperparameter.
+Here, $\alpha\in[0,1]$ is a hyperparameter.
 Model ICEWS18 ICEWS14 ICEWS05-15
@@
 RLGNet 29.90 20.18 33.64 49.08 39.06 29.34 42.03 58.12 40.83 30.06 45.91 61.93
-Table 1: Performance (in percentage) on on ICEWS18, ICESW14, and ICEWS05-15.
+Table 1: Performance (in percentage) on ICEWS18, ICEWS14, and ICEWS05-15.
 Model WIKI YAGO GDELT
@@
 5 Experiments
-5.1 Setup
 Dataset $||\mathcal{E}|$$|\mathcal{R}|$ Train Valid Test Time gap
-ICE18 23,033 256 373,018 45,995 49,545 24 hours
-ICE14 7,128 230 63,685 13,823 13,222 24 hours
-ICE15 10,488 251 368,868 46,302 46,159 24 hours
+ICEWS18 23,033 256 373,018 45,995 49,545 24 hours
+ICEWS14 7,128 230 63,685 13,823 13,222 24 hours
+ICEWS05-15 10,488 251 368,868 46,302 46,159 24 hours
 WIKI 12,554 24 539,286 67,538 63,110 1 year
@@
 Table 3: Statistics of the datasets.
-Datasets.
-We use six TKG datasets to evaluate the model’s effectiveness in entity prediction task, including ICEWS14(?), ICEWS18(?), and ICEWS05-15(?) from the Integrated Crisis Early Warning System (ICEWS) and the event-driven GDELT(?) dataset.
+5.1 Setup
+Datasets
+We use six TKG datasets to evaluate the model’s effectiveness in the entity prediction task, including ICEWS14 (?), ICEWS18 (?), and ICEWS05-15 (?) from the Integrated Crisis Early Warning System (ICEWS), and the event-driven GDELT (?) dataset.
 Public datasets WIKI (?) and YAGO (?) are also included.
 All datasets are time-partitioned into Training (80%), Validation (10%), and Test (10%).
-More details on datasets are in Table 3.
-Evaluation Metrics.
-To assess TKG reasoning performance, we used Mean Reciprocal Rank (MRR) and Hits@k metrics.
-MRR calculates average inverse rankings for actual entities across all queries, while Hits@k denotes the proportion of real entities appearing within the top k rankings.
+More details on the datasets are provided in Table 3.
+Evaluation Metrics
+To assess TKG reasoning performance, we use Mean Reciprocal Rank (MRR) and Hits@k metrics.
+MRR calculates the average inverse rankings for actual entities across all queries, while Hits@k denotes the proportion of real entities appearing within the top k rankings.
+Therefore, higher values of MRR and Hits@k indicate better model performance.
 Previous studies have indicated that traditional filtering settings are flawed (?).
 Thus, we report the experimental results after using the time-aware filtering settings.
-Baselines.
-RLGNet is compared with six baseline models on single-step reasoning task, including xERTE(?), RE-GCN(?), TITer(?), TiRGN(?), CEN(?), and RETIA(?).
-Since some models are not designed for multi-step reasoning task, we have selected a few models to report their performance on multi-step reasoning task, including CyGNet(?), RE-GCN(?), TiRGN(?), and CENET(?).
-Implementation Details.
-For all datasets, we set the vector dimension $d$ to 200, and the time vector dimension $d_{2}$ in the local history encoder to 48.
-The $top_{k}$ and $top_{k}^{all}$ are set to 20 and 200.
+Baselines
+RLGNet is compared with six baseline models on the single-step reasoning task, including xERTE (?), RE-GCN (?), TITer (?), TiRGN (?), CEN (?), and RETIA (?).
+Since some models are not designed for the multi-step reasoning task, we have selected a few models to report their performance on the multi-step reasoning task, including CyGNet (?), RE-GCN (?), TiRGN (?), and CENET (?).
+Implementation Details
+For all datasets, we set the embedding dimension to 200, and the dimension of $\textbf{V}_{1}^{L}$ in the Local History Module to 48.
+The $top_{k}$ and $top_{k}^{all}$ are set to 20 and 200, respectively.
 The number of GCN layers $\omega$ is set to 1.
@@
 The hyperparameter $\alpha$ is set to 0.8 in ICEWS, 0.9 in YAGO and WIKI, and 0.1 in GDELT.
-At the same time, a static graph constraint similar to that of RE-GCN is also added to ICEWS.
+Additionally, a static graph constraint similar to that of RE-GCN is also added to ICEWS.
 Adam is used for parameter learning, with a learning rate set to 0.001.
-In the local history encoder, we employ StepLR to adjust the learning rate.
-In this setup, the value of gamma is set to 0.8.
-For the YAGO and WIKI datasets, the step-size value is set to 2, while for other datasets, step-size is 10.
-Please note that each encoder is trained independently, which results in them not sharing weights.
+In the Local History Module, we employ StepLR to adjust the learning rate.
+In this setup, the learning rate decay is set to 0.8.
+For the YAGO and WIKI datasets, the step-size value is set to 2, while for other datasets, the step-size is 10.
+Please note that each module is trained independently, which results in them not sharing weights.
 5.2 Results
-The reasoning results of the entity prediction task are shown in Tables 1 and 2, where RLGNet outperforms other baselines in most cases.
-On six benchmark datasets, especially on the GDELT and ICEWS14 datasets, the MRR scores of RLGNet have increased by 3.42% and 2.11%, respectively.
-The primary reason for this improvement is that RLGNet not only considers repeating, local, and global historical information, but also integrates query information, which is not simultaneously considered by other models.
-In single-step reasoning performance comparison, RLGNet demonstrates stronger performance in most cases.
-In the TKG extrapolation task, methods that consider query information usually achieve higher H@1 results, in comparison to the methods that ignore it(?).
-However, these methods may perform slightly worse on other metrics like H@3 and H@10.
-This could be because methods that don’t consider query information may introduce answers that don’t fully match the specific query, leading to inferior performance on H@1.
-This is also why, in the single-step reasoning task on the ICEWS05-15 dataset, even though TiRGN surpasses RLGNet by 0.08% in H@3 and 0.53% in H@10, RLGNet still outperforms TiRGN by 1.09% in H@1.
-In multi-step reasoning task, the model is unable to know the ground truth facts during the test period.
-Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task.
-Local historical information relies on recently occurring facts, making it potentially noisier compared to single-step reasoning task.
-This is why global historical information becomes more important in multi-step reasoning task compared to single-step reasoning.
-TiRGN and RE-GCN has not fully considered global historical information, while CyGNet and CENET have overlooked local historical information.
-In contrast, RLGNet effectively integrates repetitive, local, and global historical information based on query information.
-Therefore, it achieves the best performance in multi-step reasoning task.
-5.3 Ablation Study
-Model ICEWS18 ICWS14 ICEWS05-15 WIKI YAGO GDELT
+The reasoning results of the entity prediction task are shown in Tables 1 and 2, where RLGNet outperforms other baselines in most cases.On six benchmark datasets, RLGNet has demonstrated superior performance in both multi-step and single-step reasoning tasks in most cases, particularly on the GDELT and ICEWS14 datasets, where its MRR scores have increased by 3.42% and 2.11%, respectively.
+This improvement is primarily due to RLGNet’s hybrid architecture and ensemble learning design.
+Each independent module considers historical information from different scales and constrains the answer range using candidate entities.
+The hybrid architecture design enhances the model’s performance in multi-task scenarios, ultimately leading to more accurate predictions.
+Compared to various single architecture models such as RE-GCN and TiRGCN, RLGNet exhibits stronger generalization capabilities in both multi-step and single-step reasoning.
+Moreover, compared to models that do not consider multiple scales, such as CEN and CyGNet, RLGNet shows superior performance.
+However, it is worth noting that in the single-step reasoning task on the ICEWS05-15 dataset, RLGNet’s H@3 and H@10 metrics are 0.08% and 0.53% lower than those of TiRGCN, respectively, but its H@1 metric is 1.09% higher than that of TiRGCN.
+We believe this phenomenon is because, in TKG extrapolation tasks, methods that consider candidate entities typically achieve higher H@1 results compared to those that ignore candidate entities.
+However, these methods may perform slightly worse on other metrics such as H@3 and H@10.
+This is because introducing candidate entities constrains the answer range, but candidate entities are sampled based on past facts, making the model more inclined towards past patterns.
+As a result, the model may overlook new emerging patterns or changes when predicting the future.
+Therefore, although candidate entities can improve the accuracy of the H@1 metric, they may lack diversity and breadth.
+Model ICEWS18 ICEWS14 ICEWS05-15 WIKI YAGO GDELT
 Single Multi Single Multi Single Multi Single Multi Single Multi Single Multi
@@
 Glo + Loc 34.52 29.13 45.56 38.42 50.23 39.97 81.75 63.88 87.38 78.39 24.73 20.72
-Glo + Rep 33.19 29.56 43.74 38.64 46.57 40.14 66.75 56.98 68.85 59.29 25.11 20.74
+Glo + Rep 33.19 29.56 43.74 38.64 46.57 40.14 66.75 56.98 78.85 69.29 25.11 20.74
 Loc + Rep 34.93 29.39 45.91 38.41 50.75 40.21 82.35 64.33 89.65 79.50 23.27 20.41
@@
 Table 4: The MRR (in percentage) results of the ablation studies.
-$top_{k}$ ICE18 ICE14 ICE15 WIKI YAGO GDELT 5 34.49 51.94 42.77 26.69 84.05 77.67 10 40.29 58.26 47.10 34.26 90.76 85.37 20 44.63 62.90 49.82 42.35 92.63 86.96 30 46.58 64.80 50.69 47.05 92.73 87.03 100 49.76 67.59 51.75 58.41 92.73 87.04 $\infty$ 50.42 68.38 51.85 64.93 92.73 87.04
-Table 5: The proportion of repeating facts when $top_{k}$ is set to different values.
-To validate the impact of different modules, we conducted ablation experiments on single-step and multi-step reasoning tasks on six benchmark datasets.
-The results are shown in Tables 4.
-Our ablation studies indicated that the local history encoder (Loc) and the global history encoder (Glo) do not always improve results.
-Specifically, the use of Glo and Loc modules led to a decrease in performance in single-step reasoning tasks in the ICEWS05-15 and GDELT datasets.
-However, considering that their improvement in other tasks far outweighs the decline in these two subtasks, we thus believe Glo and Loc are beneficial.
-On the other hand, the Repetitive History Encoder (Rep) can improve the final performance across various tasks, suggesting that events usually exhibit a degree of repetition.
-As shown in Table 5, the proportion of repetitive events in the dataset well corroborates this point.
-Overall, while Loc and Glo may not perform ideally in some specific tasks, these three historical information encoding methods can all contribute to better prediction results.
+Single and Multi represent single-step and multi-step reasoning, respectively.
+Since Rep can only predict repeating facts, we did not report the MRR score for Rep.
+$top_{k}$ ICEWS18 ICEWS14 ICEWS05-15 WIKI YAGO GDELT
+5 34.49 51.94 42.77 26.69 84.05 77.67
+10 40.29 58.26 47.10 34.26 90.76 85.37
+20 44.63 62.90 49.82 42.35 92.63 86.96
+30 46.58 64.80 50.69 47.05 92.73 87.03
+100 49.76 67.59 51.75 58.41 92.73 87.04
+$\infty$ 50.42 68.38 51.85 64.93 92.73 87.04
+Table 5: The proportion (in percentage) of repeating facts when $top_{k}$ is set to different values.
+5.3 Ablation Study
+To validate the impact of different modules, we conducted ablation experiments on single-step and multi-step reasoning tasks across six benchmark datasets.
+The results are shown in Table 4.
+Our ablation experiments studied the impact of the Repeating History Module (Rep), Local History Module (Loc), and Global History Module (Glo) on performance.
+From the hybrid architecture perspective, the Loc based on the RNN architecture and the Glo based on the MLP architecture can complement each other in capability.
+For instance, Glo+Loc always outperforms the individual Loc or Glo modules.
+From the multi-scale perspective, information at different scales is generally beneficial to the model’s performance.
+Analyzing from the perspective of repeating history information scale, we found that adding Rep to Loc, Glo, or Loc+Glo always improves the model’s performance.
+We believe the enhancement brought by repeating history information is due to the significant proportion of repeating facts in TKG datasets, which is evident from Table 5, where the proportion of repeating facts in the YAGO dataset reaches as high as 92.73% when $top_{k}$ is set to $\infty$.
+Regarding global and local history information scales, their performance varies due to differences between datasets (specific details can be found in Table 2).
+Since local history information only considers facts within a unit time period, the richness of facts occurring within this period determines the performance of Loc.
+For datasets with large time spans, such as the WIKI and YAGO datasets where the unit time is one year, all facts occurring within the same year are included in one time unit.
+Therefore, each unit time contains more and richer facts, which is why Loc performs better.
+However, for datasets with smaller time spans, such as the GDELT dataset where the unit time is 15 minutes, Glo performs better because it can better grasp the characteristics of the entire dataset.
+Overall, although in single-step reasoning tasks on the ICEWS05-15 and GDELT datasets, using Glo+Loc+Rep resulted in lower performance than Loc+Rep and Glo+Rep, the improvements in other tasks far exceeded the declines in these two sub-tasks.
+Therefore, we believe that the hybrid multi-scale architecture is effective in TKG reasoning tasks.
+Figure 3: The impact of $\alpha$ on MRR (in percentage) results.
 5.4 Hyperparameter Analysis
-Figure 3: The impact of $\alpha$ on MRR (in percentage) results.
-Effect of $\alpha$ Values in RLGNet.
-When processing multi-step and single-step reasoning tasks, global and local encoders contribute differently to the answers, as shown in Figure 3.
-In single-step reasoning task, we find that on the ICEWS dataset, the performance is best when $\alpha$ is between 0.6 and 0.8.
+We investigated the impact of the values of $\alpha$, $\omega$, and $top_{k}$ on the performance of RLGNet.
+The contributions of the Global and Local History Modules to the answers vary in single-step and multi-step reasoning tasks, which can be observed from the value of $\alpha$, as shown in Figure 3.
+In single-step reasoning tasks, we found that on the ICEWS dataset, the performance is optimal when $\alpha$ is between 0.6 and 0.8.
 On the WIKI and YAGO datasets, the performance peaks when $\alpha$ is between 0.8 and 1.
-On the GDELT dataset, the optimum range for $\alpha$ is between 0 and 0.2.
-Therefore, we set $\alpha$ to 0.8 and 0.1 in ICEWS and GDELT respectively, and to 0.9 in both WIKI and YAGO.
-In multi-step reasoning task, in order to achieve optimal performance, the value of $\alpha$ is usually smaller than the value of $\alpha$ in single-step inference tasks.
-This is because the local history encoder can more easily capture the sequential pattern of facts when the baseline truth is known.
-It’s noteworthy that in the GDELT dataset, the contribution of global historical information always exceeds that of local historical information.
-This is due to the GDELT dataset containing more abstract entity concepts, making it more difficult for the model to capture the sequential patterns of facts(?).
+On the GDELT dataset, the optimal range for $\alpha$ is between 0 and 0.2.
+Therefore, we set $\alpha$ to 0.8 and 0.1 for ICEWS and GDELT, respectively, and 0.9 for WIKI and YAGO.
+In multi-step reasoning tasks, the value of $\alpha$ usually needs to be smaller than in single-step reasoning tasks to achieve optimal performance.
+This indicates that the Global History Module based on MLP is more important in multi-step reasoning, whereas the Local History Module based on RNN is more crucial in single-step reasoning.
+The experimental results also demonstrate the unique advantages of the methods based on MLP and RNN architectures, validating our hybrid architecture design’s ability to leverage the strengths of both.
+Additionally, we studied the impact of $\omega$ and $top_{k}$ values on the performance of RLGNet on the ICEWS14 and YAGO datasets, as shown in Figure 4.
+As $\omega$ increases, the performance of the model on ICEWS14 fluctuates slightly, while the performance on the YAGO dataset rapidly declines.
+Therefore, we set $\omega$ to 1 for all datasets.
+In the YAGO dataset, setting $top_{k}$ at a higher value yields better performance; however, changes in $top_{k}$ do not significantly affect the performance on the ICEWS14 dataset.
+Therefore, we set $top_{k}$ for all datasets to 20.
+5.5 Effectiveness of Ensemble Learning Strategy
+To demonstrate the effectiveness of the ensemble learning strategy, we compared the performance of different modules using ensemble learning and joint learning strategies on the YAGO and ICEWS14 datasets, as shown in Table 6.
+In joint learning, all modules share the embeddings of entities and relationships, with other settings being the same as those in the ensemble learning strategy.
+We found that the performance of any module combination using the ensemble learning strategy was generally superior to that of the joint learning strategy.
+This is because, in TKG, the patterns of facts change over time, making it inevitable that there will be noisy data in the historical information.
+Although joint learning, through weight sharing, allows the model to perform better during the training phase, it also leads to fitting the noisy data, thereby impairing the model’s ability to predict future events.
+Therefore, we believe that the ensemble learning strategy is an effective approach for inference tasks in TKGs.
 Figure 4: The impact of $\omega$ and $top_{k}$ on MRR(in percentage) results in ICEWS14 and YAGO.
-Effect of $\omega$ and $tok_{k}$ Values in RLGNet.
-Figure 4 shows the impact of $\omega$ and $top_{k}$ on the MRR results in the ICEWS14 and YAGO datasets.
-As $\omega$ increases, the performance of the model on ICEWS14 fluctuates slightly, while the performance on the YAGO dataset rapidly declines, so we set $\omega$ to 1 in all datasets.
-In the YAGO dataset, setting $top_{k}$ at a higher value yields better performance however, changes in $top_{k}$ do not significantly affect the performance on the ICEWS14 dataset.
-Therefore, we set $top_{k}$ for all datasets to 20.
+Model ICEWS14 YAGO
+Ensem Joint Ensem Joint
+Glo+Loc 45.56 44.98 87.38 87.38
+Loc+Rep 45.91 45.10 89.65 88.95
+Glo+Rep 43.74 42.77 78.85 82.52
+Glo+Loc+Rep 46.15 45.40 89.69 88.55
+Table 6: The MRR (in percentage) results of the ablation studies. Ensem and Joint represent ensemble learning and joint learning, respectively.
 6 Conclusion
 In this paper, we propose RLGNet for TKG reasoning.
-RLGNet captures the characteristics of corresponding facts by using repeating, local, and global historical encoders respectively.
-In the future, finding a superior strategy to balance the contributions of global and local historical information is a worthwhile research problem.
-Experiments on six benchmark datasets prove that RLGNet outperforms existing models in most extrapolation tasks.
-7 Limitations
-In this paper, we have proposed a TKG prediction model based on three encoders. However, our approach has certain limitations:
-  • 1.
-    Due to the constraints of computational resources, the three encoders need to train independently.
-This method prevents them from sharing weights among themselves, potentially affecting the predictive performance of the model.
-  • 2.
-    In multi-step and single-step reasoning tasks, the contributions of global and local historical information should indeed be different.
-We have adopted a weighted average strategy to alleviate this issue to some extent, but choosing a specific hyperparameter $\alpha$ cannot achieve optimal performance on both multi-step and single-step reasoning tasks at the same time.
-Acknowledgments
+RLGNet is a model based on an ensemble learning strategy with a hybrid multi-scale architecture.
+By using Repeating, Local, and Global modules, RLGNet can capture historical information at different scales.
+To achieve complementary capabilities between modules and improve the model’s performance in single-hop and multi-hop reasoning tasks, RLGNet adopts a hybrid architecture design: the Local and Global modules utilize RNN and MLP architectures, respectively, and their scores are combined through weighted summation.
+We also employ an ensemble learning strategy to further enhance the model’s performance and generalization ability.
+Experiments on six benchmark datasets demonstrate that RLGNet outperforms existing models in most extrapolation tasks.
+7 Acknowledgments
 This work was supported by the National Natural Science Foundation of China (No.61866008).